# Duet: Agentic Design Understanding via Experimentation and Testing

### v1

@@ I Introduction
Let’s consider a simple example. Imagine we task an AI agent with describing a design with a reasonably complex finite state machine—for example, an implementation of $\text{I}^{2}\text{C}$. Specifically, let’s consider how an AI agent describes the $\text{I}^{2}\text{C}$ TX path clock stretching feature when it is given access to the RTL, but is not given any experimental tools. Below, we present a snippet of the output produced by such an agent on the left:
@@
Figure 1: Comparing illustrative snippets of feature-descriptions produced by a Baseline LLM flow (left) and the same flow enhanced with Duet (right), along with snippets of Duet’s experimentation and failure analysis (center).
As we will soon see, this description is missing a key detail. Imagine we use this description as a starting point for a downstream task, such as having an AI agent verify the TX path clock stretching feature. The agent will try and fail to verify the feature based on the description alone.
Now imagine we use Duet to enhance the process and produce deeper design understanding by giving the AI agent access to experimentation tools and prompting it with experimentation strategies when it generates its description. The agent decides to test its initial description using a simulation experiment. It generates a simulation testbench and calls the provided simulation tool, resulting in the log shown in the middle column above. These logs are sent back to the agent. As we can see from the log, the agent encodes its hypotheses in the testbench itself; in this case, the agent’s hypothesis (that its inputs would produce TX clock stretching behavior) proved to be false. The agent analyzes this result, as shown in the second snippet in the middle column. The agent uses the analysis to fix the testbench in later iterations, and successfully exercises the feature (not shown).
@@
Duet is a very general tactic that can be applied in any AI workflow operating over code to which experimentation tools can be applied. For example, Duet could be used in documentation generation to produce deeper insights into specific features of the design, or it could be used in formal verification to examine counterexamples to properties which the agent is struggling to prove. Though it is out of the scope of this paper (and this venue), it is likely even useful outside of hardware and EDA applications.
@@
The rest of this paper proceeds as follows. In section II, we describe the implementation of Duet. In section III, we evaluate Duet against a baseline AI workflow the task of formal verification. In section IV, we present related work. In section V we conclude.
@@ II-B Tools
The counterexample replication tool is itself an AI agent, which takes a formal property and a counterexample found by a formal verification tool and converts the counterexample into a SystemVerilog simulation testbench which replicates the counterexample. Each call to the counterexample replication tool generally involves 3–6 calls to the simulation tool, in which the agent attempts to produce a working replication. As we see in section III, this proves to be a very valuable tool.
@@ II-C Pitfalls
We have seen a few pitfalls when implementing an experimentation agent. As with all agents, experimentation agents are good at finding ways to “cheat”. For example, if not properly constrained, the counterexample replication tool will simply force internal signals of the design such that the counterexample is trivially replicated, rather than replicating it using the module inputs. Strict prompting fixes this issue. The experimentation agent can also be too lax when gathering evidence for a claim, or in the worst case, can even hallucinate results; simply asking the agent to thoroughly cite experiment logs for each of its claims can mitigate this issue.
@@ III-A Experimental Setup
    • Formal tool: runs a formal verification testbench and returns the proof status of asserts, coverpoints, and assumes.
    • Simulation tool: runs a simulation testbench and returns the logs to the agent.
    • Counterexample replication tool: a special, more directed case of the simulation tool which uses simulation to reproduce a Jasper counterexample.
@@
We run the two workflows on the same design. We use the same design summary and verification plan (steps 1 and 2 of Baseline) for both flows, to ensure we can compare the outputs property-by-property. Each flow is only given a single attempt to prove each property. We use GPT-5 for all LLM queries. We use Jasper as our formal verification tool and Verilator as our simulator.
@@ III-B Results and Analysis
In Table I, we classify a property as having been improved if Duet was able to prove a previously unproven or vacuous assertion in the Baseline. Two of the improved cases (1 and 6) were straightforward instances of the baseline flow failing to prove a property, while the Duet-enhanced flow successfully proved the property. Perhaps more interesting were the the other two cases. Property 7 was proven but vacuous in the baseline. However, via experimentation, the Duet-enabled flow was able to find a non-vacuous refinement of the property and prove it. Property 2 was unproven by the baseline, but the Duet-enabled flow successfully refined it. While not successfully proving the property, experimentation still uncovered a path towards progress (hence the $\sim$).
@@ IV-B LLM Agents for Hardware Verification
In simulation-based verification, AutoBench [20] demonstrated that LLMs can generate self-checking testbenches from RTL module descriptions using a hybrid prompt-template framework. The model achieved a 57% improvement in pass rate over baseline LLM models, but relied heavily on handcrafted scaffolding and lacked temporal reasoning. [21] extract a structured representations of design specification, which are used to prompt an LLM to obtain a set of assertions. These assertions are then refined by iteratively prompting the LLM to fix issues identified by a simulator. LAAG-RV [22] follows a similar protocol, but additionally utilizes a ‘signal-synchronization’ step to refine assertions by prompting an LLM to align the assertions with the design’s RTL code.
@@
These works confirm that while LLMs can generate hardware artifacts, domain-specific validation and feedback remain critical. Overall, these systems generally treat the LLM as a one-shot generator, not as an interactive agent capable of experimentation. Moreover, obtaining a deep understanding of the DUT is crucial for reliable verification, but limitations including limited-context windows of LLMs has prevented design-understanding techniques that can scale up with the complexity of designs [19, 18]. Approaches such as VerilogReader [34] use LLMs to enhance DUT code with natural language descriptions to provide greater context of intended functionalities in downstream LLM-tasks. Similarly, RTLExplain [35] generates top-level module summaries via a bottom-up, data-dependency-aware method: extracting a module-dependency graph, producing summaries in reversed topological order, and compiling them into a project-level knowledge base. This achieved 37% accuracy with a RAG agent, compared to 27% with conventional RAG. Once again, the design-understanding obtained through such methods is restricted by their exclusive dependence on static RTL and design specifications, whereas Duet takes advantage of powerful hardware-specific tools that allow AI agents to iteratively verify and refine their design-understanding.
@@ V Conclusion
We present Duet, a general methodology to develop comprehensive design understanding capabilities through iterative experimentation. We demonstrate how experimentation with Duet produces deeper design understanding that has concrete, positive effects on downstream tasks—in our case, formal verification.

### v2

@@ I Introduction
Let’s consider a simple example. Imagine we task an AI agent with describing a design with a reasonably complex finite state machine—for example, an implementation of $\text{I}^{2}\text{C}$. Specifically, let’s consider how an AI agent describes the $\text{I}^{2}\text{C}$ TX path clock stretching feature when it is given access to the RTL, but is not given any experimental tools. Below, in the left column of fig. 1, we present a snippet of the output produced by such an agent:
@@
Figure 1: Comparing illustrative design descriptions produced by a baseline LLM flow (left) and the same flow enhanced with Duet (right), along with snippets of Duet’s experimentation and failure analysis (center).
As we will soon see, this description is missing a key detail. Imagine we use this description as a starting point for a downstream task, such as having an AI agent verify the TX path clock stretching feature. The agent will try and fail to verify the feature based on the description alone due to a lack of detail.
Now imagine we use Duet to enhance the process and produce deeper design understanding by giving the AI agent access to experimentation tools and prompting it with experimentation strategies when it generates its description. The agent decides to test its initial description using a simulation experiment. It generates a simulation testbench and calls the provided simulation tool, resulting in the log shown in the middle column above. These logs are sent back to the agent. As we can see from the log, the agent encodes its hypotheses in the testbench itself; in this case, the agent’s hypothesis (that its inputs would produce TX clock stretching behavior) proved to be false. The agent analyzes this result, as shown in the second snippet in the middle column. The agent uses the analysis to fix the testbench in later iterations, and successfully exercises the feature (not shown). This results in the improved feature description in the right column of fig. 1
@@
Duet is a very general tactic that can be applied in any AI workflow operating over code to which experimentation tools can be applied. For example, Duet could be used in documentation generation to produce deeper insights into specific features of the design, or it could be used in formal verification to examine counterexamples to properties which the agent is struggling to prove. Though it is out of the scope of this paper, it is likely even useful outside of hardware and EDA applications.
@@
The rest of this paper proceeds as follows. In section II, we describe the implementation of Duet. In section III, we evaluate Duet against a baseline AI workflow on the task of formal verification. In section IV, we present related work. In section V we conclude.
@@ II-B Tools
The counterexample replication tool is itself an AI agent, which takes a formal property and a counterexample found by a formal verification tool and converts the counterexample into a SystemVerilog simulation testbench which replicates the counterexample. Each call to the counterexample replication tool generally involves 1–5 calls to the simulation tool, in which the agent attempts to produce a working replication. As we see in section III, this proves to be a very valuable tool as the trial-and-error process of counterexample replication seems to produce deeper design understanding.
@@ II-C Pitfalls
We have seen a few pitfalls when implementing an experimentation agent. As with all agents, experimentation agents are good at finding ways to “cheat”. For example, if not properly constrained, the counterexample replication tool will simply force internal signals of the design such that the counterexample is trivially replicated, rather than replicating it using the module inputs. Strict prompting fixes this issue. The experimentation agent can also be too lax when gathering evidence for a claim, or in the worst case, can even hallucinate results; simply asking the agent to thoroughly cite experiment logs for each of its claims can mitigate this issue. All of these observations are a reminder that it is all too easy to think of modern LLM-based AI agents as “human-like” intelligence. While they are incredibly powerful and useful, they also fail in ways most human experts never would.
@@ III-A Experimental Setup
    • Formal tool: runs a formal verification testbench with Jasper and returns the proof status of asserts, coverpoints, and assumes.
    • Simulation tool: runs a simulation testbench with Verilator and returns the logs to the agent.
    • Counterexample replication tool: a special, more directed case of the simulation tool which uses simulation via Verilator to reproduce a Jasper counterexample.
@@
We run the two workflows on the same design: a simple round-robin arbiter. We use the same design summary and verification plan (steps 1 and 2 of Baseline) for both flows, to ensure we can compare the outputs property-by-property. We arbitrarily limit the verification plan to a total of ten properties. Each flow is only given a single attempt to prove each property. We use GPT-5.0 for all LLM queries. We use Jasper as our formal verification tool and Verilator as our simulator.
@@ III-B Results and Analysis
In Table I, we classify a property as having been improved if Duet was able to prove a previously unproven or vacuous assertion in the Baseline. Two of the improved cases (1 and 6) were straightforward instances of the baseline flow failing to prove a property, while the Duet-enhanced flow successfully proved the property. Perhaps more interesting were the other two cases. Property 7 was proven but vacuous in the baseline. However, via experimentation, the Duet-enabled flow was able to find a non-vacuous refinement of the property and prove it. Property 2 was unproven by the baseline, but the Duet-enabled flow successfully refined it. While not successfully proving the property, experimentation still uncovered a path towards progress (hence the $\sim$).
@@
The evaluation also highlights where Duet can be improved. For example, despite careful prompting, there were still some cases where the agent found ways to “cheat”—for example, replicating counterexamples by simply forcing the outputs to the expected values, instead of only manipulating the inputs. Interestingly, though, some of these cases still seemed to help the agent with its understanding of the design, suggesting that simply reformatting the counterexamples into a more comprehensible format is itself useful.
Case Study: Property 1
To peek inside the Duet flow, we will look at flow artifacts for Property 1. Both the baseline and Duet-enabled flows begin by an RTL-only analysis of the design, in which the agent reads the design files and writes out a design summary, which we show in truncated form below:
Design summary (as implemented):
- Module and parameters
  - Top: arbiter
  - Parameters:
    - NUM_PORTS: number of request/grant ports ...
  - Ports:
    - clk: clock, rising-edge
    - rst: synchronous reset (active high) ...
- Core state and combinational signals ...
- Sequential behavior per clock edge ...
- What to verify (suggested properties) ...
As this summary comes only from viewing the RTL, it may be incorrect. It is on the verifier agent to realize this and refine its properties accordingly.
We then begin the actual process of verification with a prompt:
Property 001: grant_subset_of_request
Goal: Prove that in every cycle, grant is a subset of request:
  (grant & ~request) == 0.
As we will soon see, this property isn’t correct as stated; however, the agent will be able to discover this and refine the property via experimentation. After some reasoning, the agent produces its first attempt at a formal testbench. We see this behavior often—even when experimentation tools are available, the agent will often attempt to achieve the goal without them first. This behavior can be controlled via prompting, but we have found it more beneficial to focus on ensuring that the agent uses the experimentation tools if and when its initial attempts fail.
The following is a truncated version of its initial attempt at a formal testbench:
module arbiter_grant_subset_sva #(...) (...);
  property p_grant_subset;
    @(posedge clk) (grant & ~request) == ’0;
  endproperty
  assert property (p_grant_subset);
endmodule
When the agent calls Jasper on this initial attempt at a formal testbench, Jasper produces counterexamples. The agent then uses the counterexample replication tool. Recall that the counterexample replication tool is itself an agent which uses multiple calls to the simulation tool to replicate a given counterexample. Below we show the truncated output of the counterexample replication tool:
module tb_arbiter_ce;
  arbiter #(...) dut (...);
  initial clk = 1’b0; always #5 clk = ~clk;
  int cyc; initial cyc = 0; always @(posedge clk) cyc <= cyc + 1;
  // Basic stimulus to steer token to index 4 and produce a grant on bit 4
  initial begin ... end
  // Create and check the counterexample right when grant[4] goes high
  always @(posedge clk) begin
    if (cyc == 2) begin
      rst     <= 1’b1; request <= 6’h22; #1;
      assert (grant[4] == 1’b1 && request[4] == 0);
    end
  end
endmodule
Though the agent doesn’t directly use the output of the counterexample replication tool, it seems that the process of replication gives the agent a deeper understanding of the design. In this case, it shows the agent that the timing behavior of the given property is more complex than originally proposed, and helps the agent refine the property. Finally, the agent is able to produce a formal testbench which passes:
module arbiter_grant_subset_sva #(...) (...);
  property p_grant_subset_prev_req;
    @(posedge clk)
      1’b1 |-> ##1 ((grant & ~\$past(request)) == ’0);
  endproperty
  assert property (p_grant_subset_prev_req);
endmodule
Compared to the Duet-enabled flow, the baseline flow is unable to discover the exact timing of the property. While it iterates a number of times on the formal testbench, it fails to pull deeper meaning from the counterexamples returned by Jasper.
@@ IV-B LLM Agents for Hardware Verification
In simulation-based verification, AutoBench [20] demonstrated that LLMs can generate self-checking testbenches from RTL module descriptions using a hybrid prompt-template framework. The model achieved a 57% improvement in pass rate over baseline LLM models, but relied heavily on handcrafted scaffolding and lacked temporal reasoning. [14] extract a structured representation of design specifications, which are used to prompt an LLM to obtain a set of assertions. These assertions are then refined by iteratively prompting the LLM to fix issues identified by a simulator. LAAG-RV [13] follows a similar protocol, but additionally utilizes a ‘signal-synchronization’ step to refine assertions by prompting an LLM to align the assertions with the design’s RTL code.
@@
These works confirm that while LLMs can generate hardware artifacts, domain-specific validation and feedback remain critical. Overall, these systems generally treat the LLM as a one-shot generator, not as an interactive agent capable of experimentation. Moreover, obtaining a deep understanding of the DUT is crucial for reliable verification, but limitations including limited-context windows of LLMs have prevented design-understanding techniques that can scale up with the complexity of designs [8, 18]. Approaches such as VerilogReader [12] use LLMs to enhance DUT code with natural language descriptions to provide greater context of intended functionalities in downstream LLM-tasks. Similarly, RTLExplain [2] generates top-level module summaries via a bottom-up, data-dependency-aware method: extracting a module-dependency graph, producing summaries in reversed topological order, and compiling them into a project-level knowledge base. This achieved 37% accuracy with a RAG agent, compared to 27% with conventional RAG. Once again, the design-understanding obtained through such methods is restricted by their exclusive dependence on static RTL and design specifications, whereas Duet takes advantage of powerful hardware-specific tools that allow AI agents to iteratively verify and refine their design-understanding.
@@ V Conclusion
We present Duet, a general methodology to develop comprehensive design understanding capabilities through iterative experimentation. We present a generally-applicable implementation of Duet which is useful in many potential AI-assisted flows. We demonstrate how experimentation with Duet produces deeper design understanding that has concrete, positive effects on downstream tasks—in our case, formal verification.